Best friends, decided to start their own business which would bring great coffee from all over the world to their local community. Together they have formed Classic Coffee Company, or CCC (assume their coffee shop is located a few miles from our college campus). Nathan is selecting and making the coffee they serve. Cody manages the front and back office; selling coffee to customers, paying employees, and running ads for the coffee shop.

Last year, the company sold 48,800 units and had the below sales and cost figures.

Sales 134,200

Variable Costs 29,280

Contribution Margin 104,920

Fixed Costs 52,000

Income before taxes 52,920

Income Tax 15,876

Net Income 37,044

The team have been in business for 2 years. With competition rising in the coffee industry, they know that another challenging and competitive year lies ahead of them. They are considering two alternatives and have hired you to help them make some important business decision
Classic Coffee Company

Alternative 1- Bakery

Facing stiff competition from other coffee and beverage shops, Classic Coffee Company (CCC) is considering expanding its products to include bakery items. CCC would now experience higher average sales amounts per customer since beverage sales may be accompanied with food sales. This would, however, increase the company’s variable costs since it would need to purchase the food items from outside vendors. The following data pertains to adding the bakery option:

  • Total sales in the number of units would increase 25% with the added bakery option.
  • The average sales price per unit would increase from the current $2.75 to $4.00 when food sales are added to the coffee shop.
  • The variable cost per unit would increase to $1.50 per unit to account for the added cost of bakery items.
  • There would be an increase in fixed costs of $30,000 for advertising to inform the public of this change.

Alternative 2 – Drive Thru

CCC is thinking about having a drive-thru installed. Having this amenity would help the coffee shop increase its customer base, allowing them to reach a greater number of customers. This added convenience would permit CCC to slightly increase the price of its average coffee and would not increase the per unit variable costs for the company. The following data pertains to adding the drive-thru option:

  • This alternative is exclusive of alternative 1
  • Total sales in the number of units would double what it was last year.
  • The construction of the drive-thru would increase fixed costs by $130,000 – which represents one-time construction and installation costs.
  • This alternative would allow CCC to increase their average sales price to $3.30 a cup of coffee and their variable cost per unit would remain the same.

Project Requirements A & B:.

  1. Compute the Profit Margin and Return on Assets for each scenario assuming average total assets currently are $300,000; average total assets for bakery would be $500,000 and average total assets for drive-thru would be $700,000. Industry averages are 30% profit margin and 5% ROA.
Answer #1

Amounts are in $

Existing Sales price = $134,200/48,80 = $2.75

Existing Variable cost per unit = $29,280/48,800 = $0.6

Existing income tax rate = 15,876/52,920 = 30%

Bakery New units 48,60 x 125% 61,000 units New sales price = $4 Vodiable cost per unit = $15. sales - 61,000 Xb4 244000 Varia

Return on Assets on Bakery

= Net Income/Total average assets

= 49,350/800,000

= 6.17%

Return on Assets on Drive through

= 148,064/1,000,000

= 14.81%

As a long term benefit, it is better to invest in Alternative 2 (Drive thru)

Note :

1.

Return on assets is taken as Net Income/Total assets. Instead we can also take EBIT/Total assets

2.

The construction cost of $130,000 is not included in fixed cost as this is one time cost and will not occur in every years income statement. It is assumed that this construction cost is included in Total average assets given below of $700,000.

3.

Alternatively if we reduce this construction cost as fixed cost then the Profit margin would have been 17.72% and Return on assets would have been 5.71%

4.

As the useful life of the construction is not given, the depreciation on such construction and tax savings due to depreciation is ignored.